OpenStudy (chasebraves):

Solve the following equation x + 5 = 2x

OpenStudy (anonymous):

x's one one side, numbers on other

OpenStudy (anonymous):

*on one

OpenStudy (anonymous):

then you add them, and again number in front of x goes on other side and you get answer

OpenStudy (anonymous):

can you solve it by yourself or do you need further help?

OpenStudy (chasebraves):

i need further help

OpenStudy (anonymous):

okay, x + 5 = 2x 5 = 2x-x ( x changes sign because it changes side) 5 = x

alones (alones):

braves did u understand stefan steps??

OpenStudy (chasebraves):

no

alones (alones):

i can show u again :) so we have this x+5=2x Step 1: Subtract 2x from both sides. x+5−2x=2x−2x −x+5=0 Step 2: Subtract 5 from both sides. −x+5−5=0−5 −x=−5 Step 3: Divide both sides by -1. −x/−1=−5/−1(divide -5 by -1you'l have to do long division and that will be your answer) x=5 Do u understand now??

OpenStudy (chasebraves):

kind of

alones (alones):

OOOH WEL STILL IT'S BETTE TO KNOW SOMETHING THEN JUST HAVE EMPTINESS IN YOur mind avout this....I'm glad you understand at least soemthing out of my steps

OpenStudy (chasebraves):

yeah